# Hybrid optimal control with mixed-integer Lagrangian methods

### v1

@@ Abstract
Models involving hybrid systems are versatile in their application, but difficult to handle and optimize efficiently due to their combinatorial nature. This work presents a method to cope with hybrid optimal control problems which, in contrast to decomposition techniques, does not require relaxing the integrality constraints. Based on the discretize-then-optimize approach, our scheme addresses mixed-integer nonlinear problems under mild assumptions. The proposed numerical algorithm builds upon the augmented Lagrangian framework, whose subproblems are handled using successive mixed-integer linearizations with trust regions. We validate the performance of the numerical routine with extensive investigations using several hybrid optimal control problems from different fields of application. Promising preliminary results are presented for a motion planning task with hysteresis and drag, a Lotka-Volterra fishing problem, and a facility location design problem.
Keywords. Hybrid dynamics, Optimal control, Mixed-integer programming, Augmented Lagrangian.
AMS MSC. 65K05, 90C06, 90C11, 90C30.
The authors are with the University of the Bundeswehr Munich, Department of Aerospace Engineering, Institute of Applied Mathematics and Scientific Computing, 85577 Neubiberg, Germany.This research is funded by dtec.bw – Digitalization and Technology Research Center of the Bundeswehr [EMERGENCY-VRD, MissionLab, MORE, SeRANIS]. dtec.bw is funded by the European Union – NextGenerationEU.
@@ 1 Introduction
We consider hybrid optimal control problems (OCP), that is, optimization problems involving possibly nonsmooth dynamics, mixed-integer states or controls, or logical constraints. Hybrid OCPs were investigated, e.g., in [19, 14] where hybrid dynamics are modeled with disjunctive programming. An important subclass of hybrid OCPs are mixed-integer OCPs with discrete-valued controls. Numerical methods often use variable time transformations, relaxation, or decomposition techniques, compare [12, 23] and [25, 16, 17, 15, 20] for important recent developments. Such approaches are often tailored to discrete-valued controls and exploit the time dependency.
In this work, we permit more general hybrid OCPs and apply a direct discretization method to transform the dynamic optimization problem into a finite dimensional mixed-integer nonlinear program (MINLP). Our focus is not on the direct discretization method itself, which exists in various versions, compare [13] for an overview, nor on taking advantage of the time structure. We investigate instead the numerical viability of an optimization scheme that handles hybrid features as they stand, namely without relaxing integrality constraints. Details about modelling and discretization of hybrid OCPs are provided in Section 4.
@@
with $f\colon\mathcal{X}\to\mathbb{R}$ and $c\colon\mathcal{X}\to\mathbb{R}^{m}$ smooth functions, $\mathcal{C}\subset\mathbb{R}^{m}$ a nonempty closed convex set (projection-friendly in practice), and $\mathcal{X}$ a nonempty closed set with a mixed-integer linear (MIL) structure [5]. In particular, we assume $\mathcal{X}$ to be described by linear inequalities and integrality constraints on some variables, that is,
@@
for some matrix $A$, vectors $b$, $x_{\textup{l}}$ and $x_{\textup{u}}$, and index set $\mathcal{I}$. The general template (P) covers hybrid OCPs discretized with arbitrary schemes, and can capture various features such as mixed-integer control inputs, nonsmooth dynamics and state jumps.
In the following, we will partition the decision variables in $x$ into a real-valued vector $u$ and an integer-valued vector $z$, according to their admissible values for $\mathcal{X}$. Furthermore, we consider the following blanket assumptions.
@@
The approach here is to adopt the AL framework to solve the constrained problem (P) through a sequence of simpler subproblems, that is, moving the nonlinear constraints onto the objective. This methodology is motivated by sequential (partially) unconstrained minimization schemes [11], including (shifted) penalty [2] and barrier (or interior point) methods [8], and by the availability of an affordable solver [5] for problems of the form (P) without nonlinear constraints $c(x)\in\mathcal{C}$. In contrast to global optimization algorithms for (P), which often require convexity of the objective and constraint functions and become heuristics otherwise [1], our method does not rely on such convexity assumptions. Therefore, we do not seek global solutions but are content with suboptimal ones as in [21, 10].
The work contributes a solution technique for (discretized) hybrid OCPs based on an affordable safeguarded AL approach with inexact subproblems. A summary of the optimality concept is described in Section 2, while algorithmic details are provided in Section 3. High-dimensional hybrid OCPs are addressed with our method in Section 4, where numerical results demonstrate its efficacy.
@@ 2 Optimality Conditions
Since we aim at developing an affordable numerical method, possibly at the price of global optimality, we are interested in local optimality notions too. However, as these are sensitive to how neighborhoods are defined, local notions can be weak and fragile in the mixed-integer setting considered here. Indeed, considering simple balls, sufficiently small neighborhoods of feasible points may contain only one integer configuration, making it locally optimal, in a weak sense. To avoid this naive declaration of local optimality, we develop some stronger conditions by constructing neighborhoods based on $\|\cdot\|_{\textup{LP}}$, which is defined as a polyhedral norm of the real-valued components only [5]. For instance, it can take on the forms
$\|x\|_{\textup{LP}}:={}$ $\sum\{|x_{i}|\,|\,i\notin\mathcal{I}\}$ for $\ell_{1}\text{-type},$
$\text{or}\quad\|x\|_{\textup{LP}}:={}$ $\max\{|x_{i}|\,|\,i\notin\mathcal{I}\}$ for -$\ell_{\infty}\text{-type}.$
Although not a norm, $\|\cdot\|_{\textup{LP}}$ induces compact neighborhoods according to $\mathbb{B}_{\textup{LP}}(\bar{x},\Delta):=\{x\in\mathbb{R}^{n}\,|\,\|x-\bar{x}\|_{\textup{LP}}\leq\Delta\}$, owing to the boundedness assumption on integer-valued variables. Now, analogously to the “unconstrained” case [5], local minimizers for (P) can be defined as follows.
Definition 2.1 (Minimizer).
A point $\bar{x}\in\mathbb{R}^{n}$ is called a local minimizer for (P) if it is feasible and there exists $\Delta>0$ such that $f(\bar{x})\leq f(x)$ for all feasible $x\in\mathbb{B}_{\textup{LP}}(\bar{x},\Delta)$. If the latter property additionally holds for all $\Delta>0$, then $\bar{x}$ is called a global minimizer for (P).
We now focus on first-order necessary optimality conditions for (P). These should be useful and practical in order to characterize and numerically detect points that are (at least) candidates for global (or local) minimizers [2, Chapter 3].
2.1 Simple constraints
@@
Recalling [5], a first-order criticality measure associated to (2) is defined for all $x\in\mathcal{X}$ and $\Delta\geq 0$ by
$\Psi_{\varphi,\mathcal{X}}(x,\Delta):=\max_{\begin{subarray}{c}w\in\mathcal{X}\\ \|w-x\|_{\textup{LP}}\leq\Delta\end{subarray}}\langle\nabla\varphi(x),x-w\rangle\geq 0.$ (3)
@@
Definition 2.2 (Criticality).
Given some $\varepsilon>0$, a point $\bar{x}\in\mathcal{X}$ is called $\varepsilon$-critical for (2) if there exists some $\Delta>0$ such that $\Psi_{\varphi,\mathcal{X}}(\bar{x},\Delta)\leq\varepsilon$. A 0-critical point is simply called critical.
The denomination criticality, in contrast to stationarity, emphasizes that the former notion is in general significantly stronger than the latter [5].
2.2 Difficult constraints
The “unconstrained” notion of Definition 2.2 is useful to characterize solutions of a problem with simple constraints. But what is a “critical point” for a problem such as (P)? We seek a stationarity characterization that resembles, at least in spirit, the so-called Karush-Kuhn-Tucker (or KKT) conditions in nonlinear programming, see e.g. [2, 6].
@@ 2.2 Difficult constraints
From the viewpoint of nonlinear programming, where local minima correspond to saddle points of the Lagrangian functions, we consider the following notion for KKT-like points of (P), building upon Definition 2.2. Related concepts are available in [8, 6, 7].
Definition 2.3 (KKT criticality).
A point $\bar{x}\in\mathbb{R}^{n}$ is called KKT critical for (P) if $\bar{x}\in\mathcal{X}$ and there exists a multiplier $y\in\mathbb{R}^{m}$ and some $\Delta>0$ such that $\Psi_{\mathcal{L}(\cdot,y),\mathcal{X}}(\bar{x},\Delta)=0\qquad\text{and}\qquad y\in\mathcal{N}_{\mathcal{C}}(c(\bar{x})).$
Notice that KKT criticality demands feasibility since the normal cone $\mathcal{N}_{\mathcal{C}}(c(\bar{x}))$ to $\mathcal{C}$ at $c(\bar{x})$ would be empty otherwise [6]. By (3), the first condition in Definition 2.3 can be rewritten as
$\min_{\begin{subarray}{c}x\in\mathcal{X}\\ \|x-\bar{x}\|_{\textup{LP}}\leq\Delta\end{subarray}}\langle\nabla f(\bar{x})+c^{\prime}(\bar{x})^{\top}y,x-\bar{x}\rangle=0$
meaning that the Lagrangian at $\bar{x}$ cannot be (locally) further minimized with respect to $x$ while maintaining MIL feasibility, in the sense of Definition 2.2. Thanks to the $\|\cdot\|_{\textup{LP}}$-localization, this condition is in general stronger than the inclusion $0\in\nabla_{x}\mathcal{L}(\bar{x},y)+\mathcal{N}^{\textup{lim}}_{\mathcal{X}}(\bar{x})$, corresponding to mere stationarity [7], where $\mathcal{N}^{\textup{lim}}_{\mathcal{X}}(\bar{x})$ denotes the limiting normal cone to $\mathcal{X}$ at $\bar{x}$ [6]. Equivalence holds when the set $\mathcal{X}$ is convex, which is not the case as soon as $\mathcal{I}\neq\emptyset$.
Although a detailed theoretical analysis is beyond the scope of this paper, it is plausible (e.g. patterning the arguments in [6, 4]) that all local minimizers of (P) are KKT critical, possibly under some constraint qualifications.
A criterion for numerical termination arises from relaxing criticality and feasibility requirements in Definition 2.3. Analogously to $\varepsilon$-criticality of Definition 2.2, we consider the following concept, akin to those in [8, 6, 4].
Definition 2.4 ($\varepsilon$-KKT criticality).
Given some $\varepsilon\geq 0$, a point $\bar{x}\in\mathbb{R}^{n}$ is called $\varepsilon$-KKT critical for (P) if $\bar{x}\in\mathcal{X}$ and there exists a multiplier $y\in\mathbb{R}^{m}$, a vector $z\in\mathcal{C}$, and some $\Delta>0$ such that $\Psi_{\mathcal{L}(\cdot,y),\mathcal{X}}(\bar{x},\Delta)\leq{}$ $\varepsilon,$ $y\in{}$ $\mathcal{N}_{\mathcal{C}}(z),$ $\text{and}~\|c(\bar{x})-z\|\leq{}$ $\varepsilon.$ A 0-KKT critical point is simply called KKT critical.
@@ 3 Algorithmic Framework
Exploiting the fact that the augmented Lagrangian (AL) approach does not rely on smoothness [2, 6], we will employ an AL scheme to seek a numerical solution for (P) under Assumption 1.1. Building upon a sequence of shifted-penalty subproblems, AL algorithms can be described with a nested structure. An outer loop generates a sequence of primal-dual estimates, adapts shifts and penalty parameters, and monitors convergence toward KKT critical points. Then, each subproblem associated to a set of parameters is addressed by an iterative procedure, effectively an inner loop.
Although the AL framework comprises several algorithms and variants [2, 6, 4, 9], all these share a core element: the successive parametric minimization of an AL function. The safeguarded AL scheme for (P) stated in Algorithm 3.1 patterns that of [2, §4.1] and proceeds with a sequence of subproblems of the form
[…]
for some given penalty parameter $\mu>0$ and multiplier estimate $\widehat{y}\in\mathbb{R}^{m}$. Herein, the simple constraints $x\in\mathcal{X}$ are kept implicit, and the AL function with $\mu>0$ and $\widehat{y}\in\mathbb{R}^{m}$ is defined by
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract
Models involving hybrid systems are versatile in their application but difficult to optimize efficiently due to their combinatorial nature. This work presents a method to cope with hybrid optimal control problems which, in contrast to decomposition techniques, does not require relaxing the integrality constraints. Based on the discretize-then-optimize approach, our scheme addresses mixed-integer nonlinear problems under mild assumptions. The proposed numerical algorithm builds upon the augmented Lagrangian framework, whose subproblems are handled using successive mixed-integer linearizations with trust regions. We validate the performance of the numerical routine with extensive investigations using hybrid optimal control problems from different fields of application. Promising preliminary results are presented for a motion planning task with hysteresis and a Lotka-Volterra fishing problem with total variation.
Keywords. Hybrid dynamics, Optimal control, Mixed-integer programming, Constrained optimization, Augmented Lagrangian.
AMS MSC. 65K0590C0690C1190C30.
@@ 1 Introduction
University of the Bundeswehr Munich, Department of Aerospace Engineering, Institute of Applied Mathematics and Scientific Computing, 85577 Neubiberg, Germany. This research is funded by dtec.bw – Digitalization and Technology Research Center of the Bundeswehr [EMERGENCY-VRD, MissionLab, MORE, SeRANIS]. dtec.bw is funded by the European Union – NextGenerationEU. Corresponding author: Alberto De Marchi.
We consider hybrid optimal control problems (OCP), that is, optimization problems involving possibly nonsmooth dynamics, mixed-integer states or controls, or logical constraints. Hybrid OCPs were investigated, e.g., by [15, 12] where hybrid dynamics are modeled with disjunctive programming. An important subclass of hybrid OCPs are mixed-integer OCPs with discrete-valued controls. Numerical methods often use variable time transformations, relaxation, or decomposition techniques; compare [10, 18, 2] for important developments. Such approaches are often tailored to discrete-valued controls and exploit the time structure.
In this work, we permit more general hybrid OCPs and apply a direct discretization method to transform the dynamic optimization problem into a finite dimensional mixed-integer nonlinear program (MINLP). Our focus is not on the direct discretization method itself, which exists in various versions, compare [11] for an overview, nor on taking advantage of the time structure. We investigate instead the numerical viability of an optimization scheme that handles hybrid features as they stand, namely without relaxing integrality constraints. Moreover, since solving these problems to global optimality is in general computationally intractable, we do not seek global solutions and are content with suboptimal, yet practical ones, as in [8, 16].
@@
with $f\colon\mathcal{X}\to\mathbb{R}$ and $c\colon\mathcal{X}\to\mathbb{R}^{m}$ smooth functions, $\mathcal{C}\subset\mathbb{R}^{m}$ a nonempty closed convex set (projection-friendly in practice), and $\mathcal{X}$ a nonempty closed set with a mixed-integer linear (MIL) structure. In particular, we assume $\mathcal{X}$ to be described by linear inequalities and integrality constraints on some variables, that is,
@@
for some matrix $A$, vectors $b$, $x_{\textup{l}}$ and $x_{\textup{u}}$, and an index set $\mathcal{I}\subseteq\{1,\ldots,n\}$. The general template (P) covers hybrid OCPs discretized with arbitrary schemes, and can capture various features such as mixed-integer control inputs, nonsmooth dynamics and state jumps.
In the following, we will partition the decision variables $x$ into a real-valued vector $u$ and an integer-valued vector $z$, according to their admissible values for $\mathcal{X}$. Furthermore, we consider the following blanket assumptions.
@@
The approach here is to adopt the AL framework to solve the constrained problem (P) through a sequence of simpler subproblems, that is, moving the nonlinear constraints onto the objective. This methodology is motivated by sequential (partially) unconstrained minimization schemes [9], including (shifted) penalty and barrier (or interior point) methods, and by the availability of an affordable solver for problems of the form (P) without nonlinear constraints [6].
The work contributes a solution technique for (discretized) hybrid OCPs based on an affordable safeguarded AL approach with inexact subproblems. A summary of the optimality concept is described in Section 2, while algorithmic details are provided in Section 3. Hybrid OCPs are addressed with our method in Section 4, where numerical results demonstrate its modelling versatility and efficacy.
@@ 2 Optimality Conditions
Since we aim at developing an affordable numerical method, possibly at the price of global optimality, we are interested in local optimality notions too. However, as these are sensitive to how neighborhoods are defined, local notions can be weak and fragile in the mixed-integer setting considered here. Indeed, considering simple balls, sufficiently small neighborhoods of feasible points may contain only one integer configuration, making it locally optimal, in a weak sense. To avoid this naive declaration of local optimality, we develop some stronger conditions by constructing neighborhoods based on a polyhedral norm applied to the real-valued components only, denoted $\|\cdot\|_{\textup{PL}}$ as in [6, §2]. For instance, it can take on the forms
[…]
Although not a norm, $\|\cdot\|_{\textup{PL}}$ induces compact neighborhoods according to
$\mathbb{B}_{\textup{PL}}(\bar{x},\Delta):=\{x\in\mathbb{R}^{n}\,|\,\|x-\bar{x}\|_{\textup{PL}}\leq\Delta\},$
owing to the boundedness of integer-valued variables by Assumption 1.1. Now, analogously to the “unconstrained” case, local minimizers for (P) can be defined as follows.
Definition 2.1.
A point $\bar{x}\in\mathbb{R}^{n}$ is called a local minimizer for (P) if it is feasible and there exists $\Delta>0$ such that $f(\bar{x})\leq f(x)$ for all feasible $x\in\mathbb{B}_{\textup{PL}}(\bar{x},\Delta)$. If the latter property additionally holds for all $\Delta>0$, then $\bar{x}$ is called a global minimizer for (P).
Notice that non-polyhedral norms for $\|\cdot\|_{\textup{PL}}$ are deliberately avoided so that, in practice, local models of (P) take the form of mixed-integer linear programs. However, this specification is not necessary to the definition of suitable neighborhoods nor to identify stronger local minimizers.
We now focus on first-order necessary optimality conditions for (P). These should be useful and practical in order to characterize and numerically detect points that are (at least) candidates for global (or local) minimizers [1, Chapter 3].
2.1 Mixed-integer polyhedral constraints
@@
Recalling [6, Def. 2.3, Prop. 2.4], a first-order criticality measure associated to (2) is defined for all $\bar{x}\in\mathcal{X}$ and $\Delta\geq 0$ by
$\Psi_{\varphi,\mathcal{X}}(\bar{x},\Delta):=\max_{\begin{subarray}{c}x\in\mathcal{X}\\ \|x-\bar{x}\|_{\textup{PL}}\leq\Delta\end{subarray}}\langle\nabla\varphi(\bar{x}),\bar{x}-x\rangle\geq 0.$ (3)
@@
Definition 2.2.
Given some $\varepsilon\geq 0$, a point $\bar{x}\in\mathcal{X}$ is said $\varepsilon$-critical for (2) if there exists some $\Delta>0$ such that $\Psi_{\varphi,\mathcal{X}}(\bar{x},\Delta)\leq\varepsilon$. A 00-critical point is simply called critical.
As discussed in [6, §2.2], criticality is a necessary optimality condition for (2) and, as such, plays a crucial role when it comes to designing an affordable numerical method for solving (2). The denomination criticality, in contrast to stationarity, emphasizes that the former notion is in general stronger than the latter.
2.2 Nonlinear constraints
The “unconstrained” notion of Definition 2.2 is useful to characterize solutions of a problem with simple constraints. But what is a “critical point” for a problem such as (P)? A stationarity characterization that resembles, at least in spirit, the classical Karush-Kuhn-Tucker (KKT) conditions in nonlinear programming has been proposed and analysed by [4, §2.2].
@@ 2.2 Nonlinear constraints
From the viewpoint of nonlinear programming, where local minima correspond to saddle points of the Lagrangian functions, we consider the following notion for KKT-like points of (P), building upon Definition 2.2.
Definition 2.3.
A point $\bar{x}\in\mathbb{R}^{n}$ is said KKT-critical for (P) if $\bar{x}\in\mathcal{X}$ and there exists a multiplier $y\in\mathbb{R}^{m}$ and some $\Delta>0$ such that $\Psi_{\mathcal{L}(\cdot,y),\mathcal{X}}(\bar{x},\Delta)=0\qquad\text{and}\qquad y\in\mathcal{N}_{\mathcal{C}}(c(\bar{x})).$
Notice that KKT criticality demands feasibility since the normal cone $\mathcal{N}_{\mathcal{C}}(c(\bar{x}))$ to $\mathcal{C}$ at $c(\bar{x})$ would be empty otherwise. By (3)–(4), the first condition in Definition 2.3 can be rewritten as
$\min_{\begin{subarray}{c}x\in\mathcal{X}\\ \|x-\bar{x}\|_{\textup{PL}}\leq\Delta\end{subarray}}\langle\nabla f(\bar{x})+c^{\prime}(\bar{x})^{\top}y,x-\bar{x}\rangle=0,$
meaning that the Lagrangian function at $\bar{x}$ cannot be (locally) further minimized with respect to $x$ while maintaining MIL feasibility, in the sense of Definition 2.2. Thanks to the $\|\cdot\|_{\textup{PL}}$-localization, this condition is in general stronger than the inclusion $0\in\nabla_{x}\mathcal{L}(\bar{x},y)+\mathcal{N}^{\textup{lim}}_{\mathcal{X}}(\bar{x})$, corresponding to mere stationarity, where $\mathcal{N}^{\textup{lim}}_{\mathcal{X}}(\bar{x})$ denotes the limiting normal cone to $\mathcal{X}$ at $\bar{x}$. Equivalence holds when the set $\mathcal{X}$ is convex, which is not the case as soon as $\mathcal{I}\neq\emptyset$.
A criterion for numerical termination arises from relaxing criticality and feasibility requirements in Definition 2.3. Analogously to $\varepsilon$-criticality of Definition 2.2, we consider the following concept, akin to [5, Def. 3.2].
Definition 2.4.
Given some $\varepsilon\geq 0$, a point $\bar{x}\in\mathbb{R}^{n}$ is said $\varepsilon$-KKT-critical for (P) if $\bar{x}\in\mathcal{X}$ and there exists a multiplier $y\in\mathbb{R}^{m}$, a vector $z\in\mathcal{C}$, and some $\Delta>0$ such that $\Psi_{\mathcal{L}(\cdot,y),\mathcal{X}}(\bar{x},\Delta)\leq{}$ $\varepsilon,$ $y\in{}$ $\mathcal{N}_{\mathcal{C}}(z),$ $\|c(\bar{x})-z\|\leq{}$ $\varepsilon.$
Consistently with Definition 2.3, a 00-KKT-critical point is KKT-critical. Possibly under some constraint qualifications (e.g., LICQ), KKT-criticality provides a necessary optimality condition for (P). However, regardless of additional regularity, each local minimizer of (P) is asymptotically KKT-critical, cf. [4, Thm 2.6].
@@ 3 Algorithmic Framework
We employ an augmented Lagrangian (AL) scheme to seek a numerical solution for (P) under Assumption 1.1. This design choice is to exploit the fact that the AL framework maintains the original problem’s structure and it does not rely on smoothness of the objective nor on convexity of the feasible set [7]. Building upon a sequence of shifted-penalty subproblems, AL algorithms can be described with a nested structure. An outer loop generates a sequence of primal-dual estimates, adapts the penalty parameter, and monitors convergence toward KKT-critical points. Each subproblem associated to a set of parameters is addressed by an iterative procedure, effectively an inner loop. Although the AL framework comprises several algorithms and variants, all these share a core element: the successive parametric minimization of an AL function. The safeguarded AL scheme for (P) stated in Algorithm 3.1 patterns that of [1, §4.1] and proceeds with a sequence of subproblems of the form
$\operatorname*{minimize}\quad$ $\mathcal{L}_{\mu}(x,\widehat{y})$ $\operatorname{over}\quad$ $x\in\mathcal{X}$ (5)
for some given penalty parameter $\mu>0$ and multiplier estimate $\widehat{y}\in\mathbb{R}^{m}$. Herein, the MIL constraints $x\in\mathcal{X}$ are kept implicit, and the AL function is defined by
$\mathcal{L}_{\mu}(x,\widehat{y}):={}$ $f(x)+\frac{1}{2\mu}\operatorname{dist}_{\mathcal{C}}^{2}\left(c(x)+\mu\widehat{y}\right)-\frac{\mu}{2}\|\widehat{y}\|^{2},$ (6)
cf. [5, §4]. Feasibility of (5) follows from $\mathcal{X}$ being nonempty whereas well-posedness is due to continuity of $\mathcal{L}_{\mu}(\cdot,\widehat{y})$ and can be guaranteed, e.g., by coercivity or (level) boundedness arguments.
@@
$\nabla_{x}\mathcal{L}_{\mu}(x,\widehat{y})={}$ $\nabla f(x)+c^{\prime}(x)^{\top}y_{\mu}(x,\widehat{y}),$ (7a)
$\nabla_{y}\mathcal{L}_{\mu}(x,\widehat{y})={}$ $c(x)-s_{\mu}(x,\widehat{y})$ (7b)
@@
$s_{\mu}(x,\widehat{y}):={}$ $\operatorname{proj}_{\mathcal{C}}(c(x)+\mu\widehat{y}),$ (8a)
$y_{\mu}(x,\widehat{y}):={}$ $\widehat{y}+\frac{c(x)-s_{\mu}(x,\widehat{y})}{\mu}.$ (8b)